OpenStudy (anonymous):

Which constants need to be added to the right and left side of the equation to make a perfect square trinomial on the left side? 8x2 - 4x - 5 = 0 8x2 - 4x = 5 8(x2− 12x)=5 8(x^2-1/2x + __/__)=5+ __/__ remember it need balanced

OpenStudy (solomonzelman):

Do you know what a "perfect square trinomial" is? (it's ok if you don't)

OpenStudy (anonymous):

honestly no, no i don't

OpenStudy (dmcdowell):

but but but its not okay. well ya know sometimes i like to cry and just wanna companion to cry with ya know @SolomonZelman and i just need help you know @SolomonZelman please dont leave me @SolomonZelman

OpenStudy (solomonzelman):

Do you know the following property? \(\large\color{#000000 }{ \displaystyle (x+a)^2=x^2+2ax+a^2 }\)

OpenStudy (anonymous):

dude i horrible at math i'm a historian so, no i have absolutely know idea

OpenStudy (dmcdowell):

please help me @SolomonZelman cry with me man cry! @SolomonZelman

OpenStudy (anonymous):

lets take a moment to cry

OpenStudy (dmcdowell):

lmao thanks bro @mccskweasy

OpenStudy (dmcdowell):

im just screwin around and thought it wouldve be fuunny to say

OpenStudy (anonymous):

lol yeah but really need help with this

OpenStudy (dmcdowell):

one sec. hey @dakid88 could you help this guy out

OpenStudy (anonymous):

thanks man

OpenStudy (dakid88):

yep

OpenStudy (dmcdowell):

no prob @mccskweasy

OpenStudy (anonymous):

im an fan lol

OpenStudy (dmcdowell):

lmao @dakid88 is smart as f*ck. hes awesome @mccskweasy

OpenStudy (anonymous):

a fan

OpenStudy (anonymous):

im really bad at math but history is my muffican jam

OpenStudy (dmcdowell):

dang, i dont like none of it. i just started taking all my classes online cause they kicked me outta all my classes so yea. this is the easiest way to get my credits by myself

OpenStudy (anonymous):

same here actually

OpenStudy (dmcdowell):

yea man. @Kalayn321 hey are you good at math

OpenStudy (dmcdowell):

mmmmm lets see. @xMissAlyCatx can you help

OpenStudy (anonymous):

brb

OpenStudy (dmcdowell):

k

OpenStudy (xmissalycatx):

have to go :(